OpenStudy (anonymous):

-11x+y=13 in slope intercept form

OpenStudy (anonymous):

Y=mx+b Given: -11x+y=13 Add 11x to both sides to isolate the Y Your equation becomes y=13+11x Commutative property- so you can switch 13 and 11x so it becomes y=11x+13

OpenStudy (anonymous):

You need the equation in the form of y=mx+b. First step: Add 11x to both sides. This isolates y, and leaves you with: y=11x+13 That is your answer. 11/1 is the slope, and it intercepts the y axis at 13.

OpenStudy (anonymous):

Yep, perfect answer from Olivia